OpenStudy (sleepyjess):

A ship is traveling around the outside rim of an island surveying a good place to land. If it is 10 miles from a lighthouse at the center of the island and 45° from due north, how far has the ship traveled from the most northern tip of the island to the nearest tenth? 0.8 miles 4.5 miles 7.9 miles 12.5 miles

OpenStudy (sleepyjess):

@nincompoop

OpenStudy (anonymous):

Wait a second.

OpenStudy (sleepyjess):

yes?

OpenStudy (anonymous):

pi R = 8 then R = 4 ÷ π= 1.2732395447351628 cm

OpenStudy (sleepyjess):

wouldn't the radius be 10 since the ship is at the edge and the lighthouse is in the middle?

OpenStudy (anonymous):

Convert it into mile.

OpenStudy (ranga):



OpenStudy (ranga):

The entire circumference of the circle makes 360 degrees at the center. Arc length Ab makes 45 degrees at the center. They are proportional. Arc length AB / 45 degrees = Circumference of circle / 360 degrees. Arc length AB = 2 * pi *r * 45 / 360

OpenStudy (ranga):

put r = 10 above and calculate AB.

OpenStudy (ranga):

This is a direct proportions problem. 360 degrees represent circumference of the circle. 45 degrees will represent how much of the circle? 45/360 * circumference of circle = 1/8 * circumference of circle = 1/8 * 2 * pi * 10 = ?

OpenStudy (sleepyjess):

7.85?

OpenStudy (sleepyjess):

so the answer would be 7.9?

OpenStudy (ranga):

Yes, 7.9 miles.

OpenStudy (anonymous):

Yaya this was correct! (Just took the quiz) tyy